# Probing non-Gaussian correlations through entanglement generation in a many-body quantum system

### v1

@@ Abstract
Here $n_{\mathbf{k}}$ is the excitation density, $\kappa_{\mathbf{k}}$ is the anomalous pairing density, and $R_{\mathbf{k},\mathbf{q}}$ the anomalous tripling density Colussi et al. (2018b, 2020). Eventually, higher order correlations are sequentially generated, restricting the range of times addressable with this truncated set Kira (2015a); Colussi et al. (2020). In the widely used Hartree-Fock-Bogoliubov (HFB) approximation Blaizot and Ripka (1985); Ahmed-Braun et al. (2022), three-body processes in the system are neglected by setting $R_{\mathbf{k},\mathbf{q}}=0$, which results in a fully integrable and Gaussian model of the dynamics, appropriate for describing early times following the quench. Upon including the non-Gaussian cumulant $R_{\mathbf{k},\mathbf{q}}$, the additional three-body scattering processes break integrability Van Regemortel et al. (2018), and significantly improve the match between model and experiment once the gas departs from the universal prethermal state van de Kraats et al. (2024a). Additionally, it was recently shown that $R_{\mathbf{k},\mathbf{q}}$ defines an order parameter for a novel Bose-Einstein condensate of correlated triples, which is particularly sensitive to the Efimov effect characteristic for strongly interacting three-body systems Musolino et al. (2022).
@@
Results.— Having established the formalism for entropy generation in the unitary Bose gas, we will now characterize this entropy using a numerical simulation of the system. We employ the model developed in Ref. van de Kraats et al. (2024a), which formulates the many-body dynamics in terms of coupled nonlinear differential equations for a set of relevant atomic and molecular cumulants, including those presented in Eq. (1) (see Supplemental Material Sup for more details). In this model the interaction is characterized by the inverse scattering length $1/a$, which we fix at zero, and the intrinsic length $R_{*}$ Petrov (2004), which determines the molecular lifetime and characterizes the Feshbach resonance as broad ($R_{*}k_{n}\ll 1$) or narrow ($R_{*}k_{n}\gg 1$) Ho et al. (2012). As we discuss in more detail later, the intrinsic length can also be used to control the validity of the thin-shell approximation. Going forward, we will express all quantities in units of the Fermi momentum $k_{n}=(6\pi^{2}n)^{1/3}$ and Fermi time $t_{n}=2m/k_{n}^{2}$, where $n$ is the gas density and $m$ the atomic mass.
Figure 2: Trajectory through $(n_{\mathbf{k}},\absolutevalue{\kappa_{\mathbf{k}}}^{2})$ phase space for the Gaussian and beyond-Gaussian models in blue and red respectively, for $R_{*}k_{n}=0.4$ and $0\leq t/t_{n}\leq 3$. The gray region is forbidden by Eq. (6) and the white region is associated with intershell entanglement. Increasingly opague arrows show the direction of time. For visibility we rescale the axes of the four panels (a)-(d) by magnification factors $r=(1,2,10,50)$.
@@
Generally, $n_{\mathbf{k}}$ will keep growing following the quench until the system enters a quasi-equilibrated state, known as the prethermal state Van Regemortel et al. (2018); Colussi et al. (2020). Here, in the Gaussian model, $n_{\mathbf{k}}$ will plateau and start to oscillate coherently with a universal Bogoliubov frequency $\omega_{\mathbf{k}}$, analyzed in detail in Ref. Colussi et al. (2020). It defines the associated characteristic healing length $\xi k_{n}\approx\sqrt{2}$ and prethermalization timescale $\tau_{\mathbf{k}}\sim 1/\omega_{\mathbf{k}}$. For $k_{s}\xi\lesssim 1$, the characteristic excitations behave as long-wavelength phonons ($\omega_{\mathbf{k}}\sim k$), and the population oscillates slowly with respect to $t_{n}$. In the beyond-Gaussian model, we find that for these low-momentum modes the pairing amplitude itself reaches a plateau value, after which it stays approximately constant, such that the state gradually travels further into the entangled region Lat . For larger shell momenta $k_{s}\xi\gtrsim 1$ we enter the free-particle regime ($\omega_{\mathbf{k}}\sim k^{2}$), where the Gaussian model displays fast oscillations along the uncertainty boundary, seen clearly in Fig. 2. Here, in the beyond-Gaussian model, ergodic three-body correlations allow the system to break free from the prethermal state Van Regemortel et al. (2018); van de Kraats et al. (2024a), such that $n_{\mathbf{k}}$ increases beyond the maximum value attained in the Gaussian approximation. In this intermediate-time regime, the trajectory of the state through phase space becomes increasingly complicated, as seen particularly in the second and third panels of Fig. 2, however we find generally that the departure from the prethermal state is associated with a rapid damping of |$\absolutevalue*{\kappa_{\mathbf{k}}}$ |.
@@
We now focus on the beyond-Gaussian model, and use Eq. (5) to calculate the second Rényi entropy ($\gamma=2$) as a function of both shell momentum and time following the quench, plotted in Fig. 3(a). As we show in the Supplemental Material Sup , our many-body model predicts that at early times $t\ll t_{*}$, $\tilde{n}_{\mathbf{k}}$ is independent of momentum and grows according to the power law $\tilde{n}_{\mathbf{k}}\sim(t/t_{*})^{6}$. Here $t_{*}=\sqrt{t_{n}\tau}$ is the geometric mean of the Fermi time $t_{n}$ and the molecular lifetime $\tau=mR_{*}/k_{n}$ Ahmed-Braun et al. (2022), and can be interpreted as the mean transition time for the processes shown in Fig. 1. It follows that, for $\gamma\geq 1$, the Rènyi entropy obeys the same early time scaling, i.e. $s_{\gamma}(\hat{\rho})\sim(t/t_{*})^{6}$, and indeed this prediction is confirmed by our numerics as shown for $\gamma=2$ in Fig. 3(b). At intermediate times, $t\gtrsim t_{*}$, this growth is gradually slowed down, and we find that while the entropy for $k_{s}\xi\lesssim 1$ still varies only weakly with $k_{s}$, the entropy for $k_{s}\xi\gtrsim 1$ now displays a rapid decrease with the shell momentum, see Fig. 3(a). This may be attributed to the universal suppression of excitations at large momenta due to the kinetic energy, which decreases the absolute value of the two point correlation functions $n_{\mathbf{k}}$ and $\kappa_{\mathbf{k}}$ Tan (2008a, b, c); Werner and Castin (2012a, b).
@@
Its important to note that we have assumed that the initial many-body state is pure, which is only an approximation for ultracold experiments. Although theoretically more involved, notions of bipartite entanglement also exist for mixed states Amico et al. (2008), motivating an extension of our method in this direction. Going forward, our framework can be applied to study other quantum gases where non-Gaussian correlations are expected to be significant, including superfluid helium Dmowski et al. (2017); Tenart et al. (2021); Bureik et al. (2025), multicomponent Fermi gases Navon et al. (2010); Mukherjee et al. (2019); Ji et al. (2022); Schumacher et al. (2023), and Bose-Fermi mixtures Patel et al. (2023); Shen et al. (2024); van de Kraats et al. (2024b); Chuang et al. (2024). Our work also creates opportunities for further exploration of quantum resources in continuous variable systems, such as the entanglement spectrum Li and Haldane (2008) and quantum Fisher information Strobel et al. (2014); Hauke et al. (2016); Benavides-Riveros et al. (2024).

### v2

@@ Abstract
Here $n_{\mathbf{k}}$ is the excitation density, $\kappa_{\mathbf{k}}$ is the anomalous pairing density, and $R_{\mathbf{k},\mathbf{q}}$ the anomalous tripling density Colussi et al. (2018b, 2020). Eventually, as the condensate is depleted and the occupation of excited modes grows, higher order correlations are sequentially generated Kira (2015a); Colussi et al. (2020). Hence, the truncation above is justified at early times, where $n_{\mathbf{k}}$ is small van de Kraats et al. (2024a). In the widely used Hartree-Fock-Bogoliubov (HFB) approximation Blaizot and Ripka (1985); Ahmed-Braun et al. (2022), three-body processes in the system are neglected by also setting $R_{\mathbf{k},\mathbf{q}}=0$, which results in a fully integrable and Gaussian model of the dynamics, appropriate for describing early times following the quench. Upon including the non-Gaussian cumulant $R_{\mathbf{k},\mathbf{q}}$, the additional three-body scattering processes break integrability Van Regemortel et al. (2018), and significantly improve the match between model and experiment once the gas departs from the universal prethermal state van de Kraats et al. (2024a). Additionally, it was recently shown that $R_{\mathbf{k},\mathbf{q}}$ defines an order parameter for a novel Bose-Einstein condensate of correlated triples, which is particularly sensitive to the Efimov effect characteristic for strongly interacting three-body systems Musolino et al. (2022).
@@
Results.— Having established the formalism for entropy generation in the unitary Bose gas, we will now characterize this entropy using a numerical simulation of the system. We employ the model developed in Ref. van de Kraats et al. (2024a), which formulates the many-body dynamics in terms of coupled nonlinear differential equations for a set of relevant atomic and molecular cumulants, including those presented in Eq. (1) (see Supplemental Material Sup for more details). In this model the interaction is characterized by the inverse scattering length $1/a$, which we fix at zero, and the intrinsic length $R_{*}$ Petrov (2004), which determines the molecular lifetime and characterizes the Feshbach resonance as broad ($R_{*}k_{n}\ll 1$) or narrow ($R_{*}k_{n}\gg 1$) Ho et al. (2012). As we discuss in more detail later, the intrinsic length can also be used to control the validity of the thin-shell approximation. Going forward, we will express all quantities in units of the Fermi momentum $k_{n}=(6\pi^{2}n)^{1/3}$ and Fermi time $t_{n}=2m/k_{n}^{2}$, where $n$ is the gas density and $m$ the atomic mass. We limit our simulations to times $t/t_{n}<3$, where the model has been shown to give good agreement with experiment van de Kraats et al. (2024a).
Figure 2: Trajectory through $(n_{\mathbf{k}},\absolutevalue{\kappa_{\mathbf{k}}}^{2})$ phase space for the Gaussian and beyond-Gaussian models in blue and red respectively, for $R_{*}k_{n}=0.4$ and $0\leq t/t_{n}\leq 3$. The gray region is forbidden by Eq. (6) and the white region is associated with intershell entanglement. Increasingly opaque arrows show the direction of time. For visibility we rescale the axes of the four panels (a)-(d) by magnification factors $r=(1,2,10,50)$.
@@
Generally, $n_{\mathbf{k}}$ will keep growing following the quench until the system enters a quasi-equilibrated state, known as the prethermal state Van Regemortel et al. (2018); Colussi et al. (2020). Here, in the Gaussian model, $n_{\mathbf{k}}$ will plateau and start to oscillate coherently with a universal Bogoliubov frequency $\omega_{\mathbf{k}}$, analyzed in detail in Ref. Colussi et al. (2020). It defines the associated characteristic healing length $\xi k_{n}\approx\sqrt{2}$ and prethermalization timescale $\tau_{\mathbf{k}}\sim 1/\omega_{\mathbf{k}}$. For $k_{s}\xi\lesssim 1$, the characteristic excitations behave as long-wavelength phonons ($\omega_{\mathbf{k}}\sim k$), and the population oscillates slowly with respect to $t_{n}$. In the beyond-Gaussian model, we find that for these low-momentum modes the pairing amplitude itself reaches a plateau value, after which it stays approximately constant, such that the state gradually travels further into the entangled region Lat . For larger shell momenta $k_{s}\xi\gtrsim 1$ we enter the free-particle regime ($\omega_{\mathbf{k}}\sim k^{2}$), where the Gaussian model displays fast oscillations along the uncertainty boundary, seen clearly in Fig. 2. Here, in the beyond-Gaussian model, ergodic three-body correlations allow the system to break free from the prethermal state Van Regemortel et al. (2018); van de Kraats et al. (2024a), such that $n_{\mathbf{k}}$ increases beyond the maximum value attained in the Gaussian approximation. In this intermediate-time regime, the trajectory of the state through phase space becomes increasingly complicated, as seen particularly in the third and fourth panels of Fig. 2, however we find generally that the departure from the prethermal state is associated with a rapid damping of |$\absolutevalue*{\kappa_{\mathbf{k}}}$ |.
@@
We now focus on the beyond-Gaussian model, and use Eq. (5) to calculate the second Rényi entropy ($\gamma=2$) as a function of both shell momentum and time following the quench, plotted in Fig. 3(a). As we show in the Supplemental Material Sup , our many-body model predicts that at early times $t\ll t_{*}$, $\tilde{n}_{\mathbf{k}}$ is independent of momentum and grows according to the power law $\tilde{n}_{\mathbf{k}}\sim(t/t_{*})^{6}$. Here $t_{*}=\sqrt{t_{n}\tau}$ is the geometric mean of the Fermi time $t_{n}$ and the molecular lifetime $\tau=mR_{*}/k_{n}$ Ahmed-Braun et al. (2022), and can be interpreted as the mean transition time for the processes shown in Fig. 1. It follows that, for $\gamma\geq 1$, the Rènyi entropy obeys the same early time scaling, i.e. $s_{\gamma}(\hat{\rho})\sim(t/t_{*})^{6}$, and indeed this prediction is confirmed by our numerics as shown for $\gamma=2$ in Fig. 3(b). At intermediate times, $t\gtrsim t_{*}$, where nonlinear effects such as Bose enhancement of the two-body interaction become significant Kira (2015a), the entropy growth is gradually slowed down, and we find that while the entropy for $k_{s}\xi\lesssim 1$ still varies only weakly with $k_{s}$, the entropy for $k_{s}\xi\gtrsim 1$ now displays a rapid decrease with the shell momentum, see Fig. 3(a). This may be attributed to the universal suppression of excitations at large momenta due to the kinetic energy, which decreases the absolute value of the two point correlation functions $n_{\mathbf{k}}$ and $\kappa_{\mathbf{k}}$ Tan (2008a, b, c); Werner and Castin (2012a, b).
@@
It is important to note that we have assumed that the initial many-body state is pure, which is only an approximation for ultracold experiments. Although theoretically more involved, notions of bipartite entanglement also exist for mixed states Amico et al. (2008), motivating an extension of our method in this direction. Going forward, our framework can be applied to study other quantum gases where non-Gaussian correlations are expected to be significant, including superfluid helium Dmowski et al. (2017); Tenart et al. (2021); Bureik et al. (2025), multicomponent Fermi gases Navon et al. (2010); Mukherjee et al. (2019); Ji et al. (2022); Schumacher et al. (2023), and Bose-Fermi mixtures Patel et al. (2023); Shen et al. (2024); van de Kraats et al. (2024b); Chuang et al. (2024). Our work also creates opportunities for further exploration of quantum resources in continuous variable systems, such as the entanglement spectrum Li and Haldane (2008) and quantum Fisher information Strobel et al. (2014); Hauke et al. (2016); Benavides-Riveros et al. (2024).